OpenStudy (imnotgoodatmath):

Solve the equation 5(2 - 3n) = -8n - 4 for n –1 0 1 2

OpenStudy (anonymous):

first you would distribute the 5 so 5*2 and 5*-3n?

OpenStudy (anonymous):

@ImNotGoodAtMath

OpenStudy (imnotgoodatmath):

@KlOwNlOvE oh sorry what does distribute mean?

OpenStudy (anonymous):

distribute is when you have a number outside of parenthesis so in this case it's a 5 you multiply everything in parenthesis by 5

OpenStudy (imnotgoodatmath):

so like 2 x 5? and 3n x 5?

OpenStudy (anonymous):

yes but the 3 is negative so the answer to 3n x 5 will be negative

OpenStudy (anonymous):

(+-)=- (--)=+ (++)=+ positive multiplied by a negative will stay negative but 2 negatives make a positive

OpenStudy (imnotgoodatmath):

just spit balling here so it would be -1?

OpenStudy (anonymous):

10-15n=-8n-4 -15n=-8n-14 -7n=-14 how do i get n by itself

OpenStudy (imnotgoodatmath):

You muliply or divide?

OpenStudy (imnotgoodatmath):

@KlOwNlOvE

OpenStudy (anonymous):

because n is being multiplied to get it over there you divide so n=? whats negative 14 divivded by negative 7

OpenStudy (imnotgoodatmath):

my caculator is being stupid hold on

OpenStudy (imnotgoodatmath):

2?

OpenStudy (anonymous):

yes do you understand how i got to that?

OpenStudy (imnotgoodatmath):

yesh

OpenStudy (anonymous):

it isn't that hard once you understand how to do it :)